# Are Images Indistinguishable to Humans Also Indistinguishable to Classifiers?

### v3

@@ Appendix B Computational cost
Appendix C Additional results
C.1 Additional results from generative adversarial network
As shown in Tab. 14, we conducted experiments using StyleGAN-XL (Sauer et al., 2022), a state-of-the-art GAN model trained on ImageNet-256. Our results demonstrate that, despite the use of a discriminator during GAN training, the classifier can still easily distinguish between real and generated images. We argue that this is because the discriminator is trained jointly with the generator. During training, the generated data seen by the discriminator comes from a continuously evolving distribution, as the generator improves with each iteration. However, when using a classifier to distinguish between real and generated distributions, the generated distribution remains fixed.
Real dataset Generative model FID Classifier Accuracy (%)
I-256 StyleGAN-XL (Sauer et al., 2022) 2.30 ResNet-50 99.69
ViT-S 99.95
ConvNeXt-T 96.85
C.2 Results of CIFAR-10
C.3 Self-supervised classifiers for Text-to-Image distribution classification
Table 15: Four-way distribution classification on text-to-image. Linear classifier on self-supervised features also achieves high accuracy in distinguishing four distributions: COCO (Lin et al., 2014), Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024), using only 5k or 10k training samples per dataset. We use ViT-B as the backbone.
@@ C.3 Self-supervised classifiers for Text-to-Image distribution classification
C.4 Visualization of crops
@@
C.5 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
We preprocess the generated images using band-pass filters as defined in Sec. 2.2, with four threshold intervals: 0-10, 10-30, 30-50, and 50-100. An example of original and processed EDM2-XXL generated images is shown in Fig. 8. As shown in Tab. 16 and Fig. 9, for EDM2-XS and EDM2-XXL, the smallest and largest models in the EDM2 family, classifiers’ accuracy approaches random guessing (around 50%) across different threshold intervals. This indicates that for models within the same diffusion model family, which share inductive biases but differ in visual quality, classifiers unable to distinguish between them based on any specific frequency band.
@@ C.5 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
Combinations Classifier Threshold Intervals Accuracy
@@ C.6 User study
Appendix D Binary classification as a measure of distribution distance
@@
Appendix E Relationship between distribution classification and FID
The Fréchet Inception Distance (FID) is a widely used metric for evaluating the quality of generative models. It relies on feature extraction networks trained on datasets such as ImageNet and assumes that the extracted feature vectors follow a multivariate Gaussian distribution. FID calculates the Fréchet distance between these Gaussian distributions to measure discrepancies between the real and generated data. Meanwhile, as noted by Kynkäänniemi et al. (2022), FID can decrease simply by aligning the histograms of top-N classifications, without necessarily improving the perceptual quality of the generated images. Additionally, recent works like Karras et al. (2023) and Tian et al. (2024) report FID scores close to those of the ImageNet validation set, suggesting limitations in FID’s sensitivity to certain distribution differences.
@@ Appendix E Relationship between distribution classification and FID
Appendix F Comparison of rectangular and circular filters
Rectangular and circular filters are common techniques in image filtering. In this paper, we chose to implement a rectangular filter following the official FreeU implementation (Si et al., 2024), due to its simplicity and computational efficiency. For comparison, Fig. 11 presents initial results using a circular mask, denoted as U-H/2 and I-256 (Ideal). The results from both implementations are similar, and we chose to use the rectangular filter for its computational efficiency.

### v4

@@ Appendix C Computational cost
Table 12: Training time of classifiers.
Real dataset Generative model FID Classifier Accuracy (%)
I-256 StyleGAN-XL [49] 2.30 ResNet-50 99.69
ViT-S 99.95
ConvNeXt-T 96.85
I-256 SiT-XL [39] 2.06 ResNet-50 99.87
@@
Table 15: Four-way distribution classification on text-to-image. Linear classifier on self-supervised features also achieves high accuracy in distinguishing four distributions: COCO [33], Pixart-$\alpha$ [8], SDXL [43], and Playground-v2.5 [32], using only 5k or 10k training samples per dataset. We use ViT-B as the backbone.
Appendix D Additional results
D.1 Additional results from other generative models
As shown in Tab. 13, we conducted experiments using StyleGAN-XL [49], a state-of-the-art GAN model trained on ImageNet-256, as well as SiT [39], a flow matching model that extends the applicability of the method. Our results demonstrate that, despite the use of a discriminator during GAN training, the classifier can still easily distinguish between real and generated images. We argue that this is because the discriminator is trained jointly with the generator. During training, the generated data seen by the discriminator comes from a continuously evolving distribution, as the generator improves with each iteration. However, when using a classifier to distinguish between real and generated distributions, the generated distribution remains fixed.
Our experiments confirm that images generated by GANs can also be readily distinguished from real images using a classifier, despite GANs’ adversarial training approach. We have not yet conducted experiments to explore this phenomenon in other generative frameworks such as Masked Image Generation models [7, 66] or Autoregressive models [54, 55], which remains an interesting direction for future work.
D.2 Results of CIFAR-10
D.3 Self-supervised classifiers for Text-to-Image distribution classification
D.4 Visualization of crops
@@
D.5 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
We preprocess the generated images using band-pass filters as defined in Sec. A.2, with four threshold intervals: 0-10, 10-30, 30-50, and 50-100. An example of original and processed EDM2-XXL generated images is shown in Fig. 8. As shown in Tab. 16 and Fig. 9, for EDM2-XS and EDM2-XXL, the smallest and largest models in the EDM2 family, classifiers’ accuracy approaches random guessing (around 50%) across different threshold intervals. This indicates that for models within the same diffusion model family, which share inductive biases but differ in visual quality, classifiers unable to distinguish between them based on any specific frequency band.
@@ D.5 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
Combinations Classifier Intervals Accuracy
@@ D.6 User study
Appendix E Binary classification as a measure of distribution distance
@@
Appendix F Relationship between distribution classification and FID
The Fréchet Inception Distance (FID) is a widely used metric for evaluating the quality of generative models. It relies on feature extraction networks trained on datasets such as ImageNet and assumes that the extracted feature vectors follow a multivariate Gaussian distribution. FID calculates the Fréchet distance between these Gaussian distributions to measure discrepancies between the real and generated data. Meanwhile, as noted by Kynkäänniemi et al. [31], FID can decrease simply by aligning the histograms of top-N classifications, without necessarily improving the perceptual quality of the generated images. Additionally, recent works like Karras et al. [28] and Tian et al. [55] report FID scores close to those of the ImageNet validation set, suggesting limitations in FID’s sensitivity to certain distribution differences.
@@ Appendix F Relationship between distribution classification and FID
Appendix G Comparison of rectangular and circular filters
Rectangular and circular filters are common techniques in image filtering. In this paper, we chose to implement a rectangular filter following the official FreeU implementation [51], due to its simplicity and computational efficiency. For comparison, Fig. 11 presents initial results using a circular mask, denoted as U-H/2 and I-256 (Ideal). The results from both implementations are similar, and we chose to use the rectangular filter for its computational efficiency.